# Fast nonparametric feature selection with error control using integrated path stability selection

### v1

@@ Acknowledgements
Tables S3 and S4 show the top ten features for each IPSS method, ranked by their median $q$-values over the 10 CV steps, when the features are miRNA and the responses are miR-150 (Table S3) and tumor purity (Table S4), respectively. Tumor purity is the proportion of cancerous cells in a sample. Observe that the $q$-values in Table S3 are much smaller than those in Table S2. This is to be expected since the signal between miRNAs, in particular, between miR-150 and certain other miRNAs, is likely much stronger than the signal between miRNAs and prognosis.
IPSSGB IPSSRF IPSSL
miR-142(0.031) miR-146a(0.016) miR-146a(0.028)
miR-146a(0.031) miR-142(0.016) miR-142(0.028)
miR-1247(0.031) miR-223(0.016) miR-1247(0.028)
miR-642(0.032) miR-642(0.016) miR-139(0.028)
miR-342(0.036) miR-155(0.016) miR-342(0.029)
miR-139(0.042) miR-1247(0.016) miR-642(0.029)
miR-1249(0.051) miR-766(0.016) miR-96(0.029)
miR-96(0.064) miR-342(0.017) miR-511-1(0.03)
miR-204(0.072) miR-1228(0.018) miR-34b(0.032)
miR-34b(0.074) miR-21(0.02) miR-184(0.033)
Table S3. MicroRNA and miR-150. Top ranked miRNAs for each method in terms of their $q$-values, shown in parentheses.
IPSSGB IPSSRF IPSSL
miR-150(0.05) miR-150(0.02) miR-150(0.05)
miR-22(0.05) miR-22(0.02) miR-22(0.05)
miR-140(0.08) miR-140(0.02) miR-142(0.05)
miR-142(0.08) miR-142(0.02) miR-140(0.05)
miR-1224(0.12) miR-146a(0.03) miR-25(0.06)
miR-103-2(0.16) miR-223(0.03) miR-1224(0.09)
miR-223(0.23) miR-214(0.04) miR-135a-2(0.1)
miR-181c(0.36) miR-132(0.04) miR-155(0.11)
miR-155(0.49) miR-642(0.05) miR-1247(0.12)
miR-25(0.52) miR-199b(0.05) miR-34c(0.13)
Table S4. MicroRNA and tumor purity. Top ranked miRNAs for each method in terms of their $q$-values, shown in parentheses.
S3.2.  RNA-seq and ovarian cancer.
In this section, we perform a similar analysis to the miRNA study in Sections 5 and S3.1, but with the 6426 genes from the RNA-seq data described in Section 4.
@@
Gene Citations Publications IPSSGB IPSSRF IPSSL
AKT2 9088 118 0.25
CD38 852 30 0.11 0.31
PPL 151 8 0.8 0.26
HOXB4 127 7 0.36
ZFHX4 105 7 0.78 0.17
SLAMF7 101 2 0.38 0.36
WNK1 96 3 0.42 0.23
SLAMF1 58 5 0.38
PARD6B 61 1 0.28 0.33
UGT8 37 1 0.17 0.41
AP2A1 37 2 0.4
PROSC 0 0 0.33 0.46 0.2
MRAP2 0 0 0.09 0.39 0.78
OR51A4 0 0 0.59 0.2
CCDC84 0 0 0.2
SH3BGR 0 0 0.07
Table S5. RNA-seq and prognosis. Genes are ordered by citations. A missing $q$-value means that gene was not selected by the corresponding version of IPSS. *A literature search revealed that the gene MRAP2 (also called C6orf117), while not directly related to ovarian cancer, is very much related to obesity, which is linked to poor prognosis and reduced survival times in cancer patients (Chan et al., 2009).
Figure S8. RNA-seq and prognosis. No method does well at predicting prognosis relative to the full dataset. This is unsurprising, given the many factors that can contribute to poor outcomes which are unlikely to be captured by RNA-seq data alone. Nevertheless, several of the genes identified by the various feature selection methods do appear to be related to ovarian cancer prognosis; see Table S5.
Figure S9. RNA-seq and AKT2. As in several of the other studies, IPSSGB achieves superior predictive performance with fewer features compared to other methods.
IPSSGB IPSSRF IPSSL
SUPT5H(0.01) SUPT5H(0.01) SUPT5H(0.02)
EGLN2(0.01) PAK4(0.01) EGLN2(0.02)
PAK4(0.01) ADCK4(0.01) ADCK4(0.02)
ADCK4(0.01) EGLN2(0.01) BLVRB(0.02)
BLVRB(0.01) ZNF574(0.01) SPSB3(0.03)
ITPKC(0.02) GSK3A(0.01) HKR1(0.03)
HKR1(0.04) SARS2(0.01) CAV2(0.04)
ZMYND17(0.14) ITPKC(0.01) MAP3K10(0.07)
U2AF2(0.19) CLPTM1(0.01) ACRC(0.08)
SPSB3(0.2) ZNF526(0.01) ITPKC(0.09)
Table S6. RNA-seq and AKT2. Top ranked genes for each method in terms of their $q$-values, shown in parentheses.

### v2

@@ Acknowledgements
miRNA Citations IPSSGB IPSSRF IPSSL1 KOGLM KORF KOL1 DeepPINK SSBoost
miR-93 43996 0.35 0.11
miR-148a 42177 0.39
miR-150 41195 0.23 0.39
miR-96 23010 0.23 0.47
miR-342 20291 0.39 0.23
miR-30d 19267 0.23 0.33
miR-301b 3224 0.35
miR-1270 2543 0.28 0.11 0.21
miR-1301 1390 0.35
miR-1-2 1220 0.35 0.21
$\geq$ 1000 8 6 4 0 0 0 0 0
$<$ 1000 0 0 0 0 0 0 0 0
Total 8 6 4 0 0 0 0 0
Table S6. MicroRNAs and prognosis (ovarian cancer). MiRNAs are ordered by citation count. A missing $q$-value indicates the miRNA was assigned a $q$-value of less than 0.5 by the corresponding method. The bottom rows report, for each method, the number of selected features with over 1000 citations, under 1000 citations, and the total number selected at the maximum target FDR of 0.5.
miRNA Citations IPSSGB IPSSRF IPSSL1 KOGLM KORF KOL1 DeepPINK SSBoost
miR-21 190920 0.08
miR-155 128839 0.32 0.08 0.09 0.20
miR-145 89265 0.24
miR-146a 57373 0.03
miR-214 57253 0.05
miR-223 53770 0.24 0.03 0.09
miR-25 43636 0.12 0.32 0.09 0.20
miR-22 42211 0.04 0.03 0.07 0.20 0.17
miR-150 41195 0.04 0.03 0.07 0.20 0.17
miR-142 39980 0.04 0.03 0.25
miR-335 33200 0.26
miR-15b 30988 0.24 0.20 0.25
miR-140 29667 0.04 0.03 0.17
miR-152 25767 0.16
$\geq$ 1000 14 25 12 0 0 7 0 5
$<$ 1000 0 3 1 0 0 1 0 0
Total 14 28 13 0 0 8 0 5
Table S7. MicroRNAs and tumor purity (ovarian cancer). MiRNAs are ordered by citation count. A missing $q$-value indicates the miRNA was assigned a $q$-value of less than 0.35 by the corresponding method. The bottom rows report, for each method, the number of selected features with over 1000 citations, under 1000 citations, and the total number selected at the maximum target FDR of 0.35.
@@
Gene Citations IPSSGB IPSSRF IPSSL1 KOGLM KORF KOL1 DeepPINK SSBoost
CD38 109641 0.13 0.14
AKT2 97889 0.13 0.30
ERBB4 61018 0.45
CCR3 25401 0.23 0.35
CD1C 24420 0.38
WTAP 22418 0.24
SHMT2 19844 0.30
AAAS 16226 0.30 0.45
PAK4 14396 0.38
SLAMF7 12027 0.24 0.14
$\geq$ 200 15 14 8 13 2 7 0 0
$<$ 200 4 1 3 4 1 0 0 0
Total 19 15 11 17 3 7 0 0
Table S8. RNA-seq and prognosis (ovarian cancer). Genes are ordered by citation count. A missing $q$-value indicates the gene was assigned a $q$-value of less than 0.5 by the corresponding method. The bottom rows report, for each method, the number of selected features with over 200 citations, under 200 citations, and the total number selected at the maximum target FDR of 0.5.
miRNA Citations IPSSGB IPSSRF IPSSL1 KOGLM KORF KOL1 DeepPINK SSBoost
miR-155 94172 0.05
miR-10b 42000 0.21 0.05 0.03 0.35 0.39
miR-148a 32559 0.13
miR-335 22423 0.08 0.05 0.50
miR-15b 21960 0.14 0.05 0.03 0.20 0.38
miR-424 21680 0.37
miR-10a 20065 0.21
miR-224 18932 0.16 0.38
miR-503 11894 0.14 0.06 0.38
let-7e 11348 0.25 0.06 0.38
$\geq$ 100 15 21 6 0 0 14 0 14
$<$ 100 5 2 3 0 0 10 0 4
Total 20 23 9 0 0 24 0 18
Table S9. MicroRNAs and prognosis (glioma). MiRNAs are ordered by citation count. A missing $q$-value indicates the miRNA was assigned a $q$-value of less than 0.5 by the corresponding method. The bottom rows report, for each method, the number of selected features with over 100 citations, under 100 citations, and the total number selected at the maximum target FDR of 0.5.
Gene Citations IPSSGB IPSSRF IPSSL1 KOGLM KORF KOL1 DeepPINK SSBoost
FOXM1 62538 0.25
WEE1 26648 0.10 0.06 0.44
IGFBP2 24482 0.08
CX3CL1 23044 0.34
TIMP1 22632 0.44
SKI 19220 0.12 0.23 0.03 0.44 0.34
CCNB1 14856 0.10
CDK9 14558 0.25
TOP2A 13820 0.03
PDPN 12929 0.23
MSN 11866 0.10
ATF2 11040 0.10
$\geq$ 500 19 18 12 0 22 24 0 0
$<$ 500 9 6 7 0 15 36 0 0
Total 28 24 19 0 37 60 0 0
Table S10. RNA-seq and prognosis (glioma). Genes are ordered by citation count. A missing $q$-value indicates the gene was assigned a $q$-value of less than 0.5 by the corresponding method. The bottom rows report, for each method, the number of selected features with over 500 citations, under 500 citations, and the total number selected at the maximum target FDR of 0.5.
S4.4  Cross-validation.
As noted in the main text, we also measure feature selection performance by implementing a 20-fold cross-validation (CV) procedure, described as follows. In each of the 20 CV steps, one group of patients is set aside (the test set), and a set of features is selected by each method using the data in the remaining groups (the training set). Next, for each method, we construct three predictive models—a linear model, a random forest model, and a gradient boosting model—using only the features selected by that method on the training data. Each model is then used to predict responses from the test set, and the smallest of the three prediction errors is recorded (we use mean squared error for regression and $1-\mathrm{accuracy}$ for classification). All three models are implemented to ensure that no method has an inherent advantage over another. For example, the features selected by IPSSL1 may be better suited to minimizing error in a linear model than those selected by IPSSGB, while those selected by IPSSGB may be better suited to minimizing error in a gradient boosting model than the ones selected by IPSSL1. The linear and random forest predictive models are implemented with scikit-learn (Pedregosa et al., 2011) and gradient boosting with XGBoost (Chen and Guestrin, 2016), always with default parameters. For continuous responses, in each CV step we subtract the mean of the training responses from all responses, training and test, and scale all responses by the empirical standard deviation of the training responses.
CV study results are shown in Figures S6S11. Each plot contains two subplots. The left subplots show the prediction error associated to the features selected by each method at the given target FDR. In all studies, we find that the three IPSS methods select features at much lower target FDRs than all of the model-X knockoffs methods. The right subplots show the prediction error as a function of the number of features selected by each method. Curves for each method in these plots are obtained by varying the target FDR between 0 and 0.5. Boruta does not have FDR control parameters and is therefore represented by a single point in these plots. In each plot, the dashed black line shows the average error when using all features in the dataset to predict the response variable. DeepPINK rarely selects any features and is therefore represented by a single point for better visibility.
Figure S6. MiRNAs and tumor purity (ovarian cancer).
Figure S7. MiRNAs and miR-150 (ovarian cancer).
Figure S8. Genes and AKT2 (ovarian cancer).
Figure S9. MiRNAs and miR-155 (glioma).
Figure S10. Genes and prognosis (glioma).
Figure S11. Genes and FOXM1 (glioma).
S5.  Sensitivity analyses
Figures S12S19 show the sensitivity of IPSSGB to the IPSS parameters discussed in Sections 2 and S1.3. Data are simulated according to the ovarian cancer RNA-seq simulation design described in Sections 3.3 and S3.2 for both regression and classification. The results for IPSSRF were similar and are therefore omitted.
Figure S12. Different choices of $C$ (regression).
Figure S13. Different choices of $C$ (classification).
Figure S14. Different choices of $\delta$ (regression).
Figure S15. Different choices of $\delta$ (classification).
Figure S16. Different choices of $B$ (regression).
Figure S17. Different choices of $B$ (classification).
Figure S18. Different choices of function (regression).
Figure S19. Different choices of function (classification).